Can you help me solve x+2/5=x+3/10
Answer by tutorcecilia(2152)

x+2/5=x+3/10 [Multiply each term by the LCD=(5)(10)=50]
(50)(x+2/5)=(50)(x+3/10)[Cancel wherever possible]
(10)(x+2)=(5)(x+3)
10x+20=5x+15
10x-5x=15-20
5x=-5
x=-1
.
check by pluggig (x=-1) back into the original equation